OpenStudy (linda):

solve for theta: sec theta = csc (2 theta - 15) please help! desperateeee

OpenStudy (watchmath):

Since sec =1/cos and csc=1/sin by flipping the equation we have cos(theta)=sin(2theta -15) Can you solve that now?

OpenStudy (linda):

i dont know how to solve it though. but i understand what you just said. could u explain the next step?

OpenStudy (watchmath):

firs is the problem \(sin(2(\theta-15))\) or \(\sin(2\theta-15)\) ?

OpenStudy (linda):

csc (2θ−15)

OpenStudy (watchmath):

Remember that \(cos \theta =\sin(90-\theta)\)

OpenStudy (watchmath):

So one of the solution can be obtain by setting \(90-\theta =2\theta -15\)

OpenStudy (linda):

ok so i got 105 = theta...is that the answerr?